# Random regular graph states are complex at almost any depth

### v1

@@ Abstract
Graph states are fundamental objects in the theory of quantum information due to their simple classical description and rich entanglement structure. They are also intimately related to IQP circuits, which have applications in quantum pseudorandomness and quantum advantage. For us, they are a toy model to understand the relation between circuit connectivity, entanglement structure and computational complexity. In the worst case, a strict dichotomy in the computational universality of such graph states appears as a function of the degree $d$ of a regular graph state [GDH+23]. In this paper, we initiate the study of the average-case complexity of simulating random graph states of varying degree when measured in random product bases and give distinct evidence that a similar complexity-theoretic dichotomy exists in the average case. Specifically, we consider random $d$-regular graph states and prove three distinct results: First, we exhibit two families of IQP circuits of depth $d$ and show that they anticoncentrate for any $2<d=o(n)$ when measured in a random $X$-$Y$-plane product basis. This implies anticoncentration for random constant-regular graph states. Second, in the regime $d=\Theta(n^{c})$ with $c\in(0,1)$, we prove that random $d$-regular graph states contain polynomially large grid graphs as induced subgraphs with high probability. This implies that they are universal resource states for measurement-based computation. Third, in the regime of high degree ($d\sim n/2$), we show that random graph states are not sufficiently entangled to be trivially classically simulable, unlike Haar random states. Proving the three results requires different techniques—the analysis of a classical statistical-mechanics model using Krawtchouck polynomials, graph theoretic analysis using the switching method, and analysis of the ranks of submatrices of random adjacency matrices, respectively.
@@ 1 Introduction
In our work, we consider graph states from the perspective of understanding quantum properties that lead to computational speedups. Graph states are prime candidates to study a specific quantum phenomenon, namely, entanglement, as it relates to computational complexity. On the one hand, measuring graph states in adaptive single-qubit bases allows the execution of arbitrary quantum computations [56, 57]. In fact, these measurements can be restricted to the $X$-$Y$ plane of the Bloch sphere [47]. On the other hand, graph states exhibit a rich multipartite entanglement structure [34] which is in fact required for quantum speedups [16]. Understanding which properties of graph states makes them hard to simulate classically can therefore yield insights into the mechanisms underlying quantum speedups.
A first step in this direction has been made in Ref. [23], where it was shown that the worst-case complexity of simulating regular graph states, i.e., graph states whose underlying graph on $n$ vertices is $d$-regular for some $0\leq d\leq n$, undergoes two sharp transitions as $d$ is increased: For $d\leq 2$ and $d\geq n-3$ simulating single-qubit measurements in any basis is classically easy, while for any other value of $d$ it is worst-case hard over the choice of $d$-regular graphs. At the very same thresholds of $d$, these graph states also undergo a transition in multipartite entanglement, providing a tight link between complexity and entanglement. This result immediately poses the question whether the graphs identified in Ref. [23] underlying intractable states are isolated in their respective regularity class, or whether the result is generic, i.e., hold for (uniformly) random $d$-regular graph states. These states form an ensemble of graph states with a well-controlled connectivity and can be efficiently sampled from [29]. The average-case hardness of random $d$-regular graph states would significantly strengthen the connection between the connectivity of a graph and the entanglement properties of the corresponding state and provide complexity-theoretic evidence that classical intractability is a generic feature of multipartite entangled states.
From a different perspective, graph states are intrinsically connected—in fact, instances of—another family of quantum states generated by commuting instantaneous quantum polynomial (IQP) circuits [63, 64]. These states, also known as phase states, have in the past received broad attention in the context of demonstrating quantum advantage. This is because it has been shown that a classical algorithm sampling from random IQP circuits would imply a collapse of the polynomial hierarchy subject to the robust average-case #P-hardness of a certain computational problem [7, 8]. IQP circuits have received revived attention recently as a means to generate quantum pseudorandomness [49, 4, 48]. They have also been used as a sandbox to study cryptographic hardness conjectures [20], as well as the power of resources like intermediate measurements [43]. Since they turn out to be natural circuits which can be implemented in a fault-tolerant way in certain stabilizer codes [55, 36], they are amenable to near-term fault-tolerant implementations [3].
Since graph states are prepared by applying controlled-phase gates to an initial |$\ket{+}^{\otimes n}$ state, the degree of the underlying graph is in close correspondence with the circuit depth required to prepare them (all $d$-regular graph states can be prepared in depth at most $d+1$). The specific IQP circuits corresponding to hexagonal and grid graphs generate universal resource states for measurement-based quantum computation in constant depth and have been shown to be classically intractable to simulate [30, 5, 35]. But it has been an open question whether or not there are ensembles of IQP circuits which are classically intractable for any depth above a constant threshold, and the best lower bound we had for such a threshold—or rather evidence for it—has been logarithmic depth [8, 36]. In particular, it is unknown, whether states which are not exploitable resources for measurement-based quantum computation (i.e., are universal) yield hard-to-simulate sampling problems. At the same time, it is unknown whether highly connected graph states are generically universal for measurement-based quantum computing and if so in what sense, or even whether they are hard to simulate classically. They might also be sufficiently entangled to be amenable to a trivial simulation algorithm [24].
@@
In this paper, we address both of these problems by considering the average-case complexity of simulating uniformly random $d$-regular graph states when measured in an arbitrary $X$-$Y$ plane product basis. Importantly, the setting we consider covers both sampling problems and measurement-based quantum computing. We give evidence for the average-case complexity of these tasks in three distinct regimes of the regularity parameter, (1) the regime of $2<d=O(1)$, (2) the regime of $d\in\Theta(n^{c})$ for any $11/2<c<$ 1, and (3) the regime of $d\sim n/2$.
@@
We first give evidence for average-case complexity in the regime of $2<d=O(1)$. To this end, we first show that the output distributions of IQP circuits on random highly connected graphs constructed from $d$ random matchings measured in a random $X$-$Y$ plane basis have the anticoncentration property [1, 31] at any depth $2<d=o(n)$. Combined with the fact that for each value of $d$ there are worst-case-hard instances in this family, this provides evidence for hardness of simulation on the same level of rigor as is known for other discrete families of IQP circuits [8, 33]. This is the first family of quantum circuits that we are aware of for which there is evidence of simulation hardness at any sublinear depth above a constant threshold, and thus a result of independent interest. We use this result to characterize the complexity of random regular graph states, noting that a random circuit from this matching model (strictly speaking a slight variation called the pairing model) at a fixed depth $d$ yields a uniformly random $d$-regular graph state with probability $O(2^{-d^{2}})$. This implies the anticoncentration property of constant-regularity graph states and thus gives evidence for their classical intractability, showing the first main result.
@@
In our second result, we show that a random $d$-regular graph with $d=\Theta(n^{c})$ for any $11/2<c<$ 1 has a grid graph of polynomial size as an induced subgraph (resulting from deleting some of the vertices). Since the grid graph is a universal resource for measurement-based quantum computation, this shows that an algorithm for sampling from the output distribution of those graph states in any local basis would imply a collapse to the polynomial hierarchy up to the average-case #P-hardness of a certain approximation problem (as well as being, in principle, universal resources for MBQC).
@@
Finally, in our third result, we show that uniformly random graph states have geometric entanglement bounded significantly away from the maximum. This makes them nontrivial and in particular not amenable to the trivial simulation algorithm of Ref. [24], giving some evidence for their average-case complexity when measured in any local basis. We believe similar results will hold for random $cn-$regular graphs with $20<c<1/$ 2.
@@ 1.1 Guide for readers
  • Anticoncentration results for IQP circuits and constant-degree regular graph states: A summary of the results and proof techniques is given in Section 1.2.1, Section 1.2.2, Section 1.3.1. The results are discussed in detail and proven in Section 3 with preliminaries in Sections 2.1, 2.2 and 2.3.
  • Universality results for random regular graph states of intermediate degree: A summary of the results and proof techniques is given in Sections 1.2.3 and 1.3.2. The results are proven in Section 4.
  • Absence of a geometric entanglement barrier for random graph states of high degree: The main results and proof techniques are summarized in Sections 1.2.4 and 1.3.3. The results are proven in Section 5 with preliminaries in Section 2.4.
Not to be missed in either case is the discussion in Section 1.4.
@@ 1.2 Results
We present results on uniformly random $d$-regular graph states covering three different ranges for the regularity parameter $d$. We deal, in order, with $2<d=O(1)$, $d=\Theta(n^{c})$ for $c\in(0,1)$, and $d\sim n/2$. Along the way we will also prove several results for different random graph models, such as the pairing and matching models (described in Section 2.2) and the uniformly random graph model. We believe these results to be of independent interest.
@@ 1.2.1 Anticoncentration of a family of random pairing and matching IQP circuits
Our first set of results deals with the average-case hardness of two families of IQP circuits of depth $d$ satisfying $d>2$ and $d=o(n)$. These IQP circuits prepare (not always regular) graph states of degree at most $d$. Average-case hardness of constant-regularity graph states will follow from those results. To give evidence for average-case hardness of these depth-$d$ circuits we prove that their outcome distribution, when measured in a random local basis in the $X$-$Y$ plane anticoncentrates. Moreover, they contain worst-case hard instances.
@@
Specifically, we consider the following two ensembles of depth-$d$ IQP circuits, which generate graph states of degree $d$. We call the first ensemble the random pairing model. This model is obtained by choosing a uniformly random matching on $n\cdot d$ vertices and identifying $d$ vertices with a qubit. A $CZ$ gate is applied to every edge in the resulting multigraph. This model (which is well studied in graph theory [74]) is motivated by the fact that, conditioning on the graph being simple, it generates a uniformly random $d$-regular graph state. We associate (simple) graph states to the multigraphs chosen in these ways by deleting double edges and self-loops. In the second model, the random matching model, a graph state is generated by applying $CZ$ gates on $d$ independent, uniformly random matchings. This model is motivated by recent experiments on universal circuits on random matchings in reconfigurable atomic quantum computers [12], where a quick convergence to simulation hardness was observed with the depth. In this model, for any constant $d$, a regular graph state is generated with constant probability. However, not all $d$-regular graph states admit a perfect matching [42] (and some admit many, making this distribution different from the uniformly random regular graphs, even when conditioning on simplicity). Thus, the regular graph states generated by perfect matchings are exactly those graph states with $dn/2$ edges which can be prepared in optimal depth $d$.
Figure 1: Exact values of $\mathbb{E}_{G,\theta}[m_{2}(G,\theta)]$ (evaluated via Equation 106) for graph states drawn from (a) the random pairing ensemble $\mathcal{G}_{p}(d,n)$ and (b) the random matching ensemble $\mathcal{G}_{m}(d,n)$ for various values of $d$ as a function of $n$. The opaque (transparent) loosely dotted lines denote the asymptotic values of $m_{2}=2(3)$ of the average second moment. Solid (dashed) lines represent odd (even) values of $d$.
We show that the output distribution of both models in a random $X$-$Y$ basis anticoncentrates for any $d$ satisfying $d>2$ and $d=o(n)$.
Consider random graph states |$\ket{G}$ on $n$ vertices drawn from (a) the random pairing model on $nd$ vertices or (b) $d$ uniformly random matchings on $n$ vertices. Then for $2<d=o(n)$
$\mathbb{E}_{G,\theta}[m_{2}(G,\theta)]=o(1)+\begin{cases}2&\text{ if $d$ is odd }\\ 3&\text{ if $d$ is even }.\end{cases}$ (1)
We also show the exact value of $\mathbb{E}_{G,\theta}[m_{2}(G,\theta)]$ for different values of $n,d$ and both models in Figure 1. We observe not only that the $d$-dependence of our results is tight, but also that the convergence to the asymptotic value occurs rapidly.
@@ 1.2.2 Anticoncentration of constant-degree regular graph states
Consider the uniform measure $\mathcal{G}_{r}(n)$ on $d$-regular graphs on $n$ vertices, and uniformly random angles $\theta$. Then, for $2<d=o(n)$,
@@
[…]
We note that some existing results about IQP circuits can be rephrased in terms of random graph state models yielding results similar to Corollary 2 for different models of random graphs. In particular, Ref. [36] shows that IQP circuits composed of $g$ $CZ$ and $Z$ gates applied uniformly at random to $n$ qubits anticoncentrate if $g=\Omega(n\log n)$ but fail to anticoncentrate for any $g=O(n)$. This directly translates to the Erdős-Rényi random graph model in which an edge is contained in the graph with probability $p=g/\binom{n}{2}$, measured in the ±$\pm X$ basis. Ref. [36] gives evidence that Erdős-Rényi random graphs are hard to simulate for $p\in\Omega(\log n/n)$ and suggests that they are not generically hard to simulate below that threshold. This is consistent with graph percolation, where famously Erdős and Rényi [19] showed that if $p<1/2n$, a random graphs largest connected component has size $O(\log n)$ which implies that it is efficiently simulatable using tensor-network techniques [16].

### v2

@@ Abstract
Graph states are fundamental objects in the theory of quantum information due to their simple classical description and rich entanglement structure. They are also intimately related to IQP circuits, which have applications in quantum pseudorandomness and quantum advantage. For us, they are a toy model to understand the relation between circuit connectivity, entanglement structure and computational complexity. In the worst case, a strict dichotomy in the computational universality of such graph states appears as a function of the degree $d$ of a regular graph state [GDH+23]. In this paper, we study the average-case complexity of simulating random graph states of varying degree when measured in random product bases and give distinct evidence that a similar complexity-theoretic dichotomy exists in the average case. Specifically, we consider random $d$-regular graph states and prove three distinct results: First, we exhibit two families of IQP circuits of depth $d$ and show that they anticoncentrate for any $2<d=o(n^{1/2})$ when measured in a random $X$-$Y$-plane product basis. This implies anticoncentration for random constant-regular graph states. Second, in the regime $d=\Theta(n^{c})$ with $c\in(0,1)$, we prove that random $d$-regular graph states contain polynomially large grid graphs as induced subgraphs with high probability. This implies that they are universal resource states for measurement-based computation. Third, in the regime of high degree ($d\sim n/2$), we show that random graph states are not sufficiently entangled to be trivially classically simulable, unlike Haar random states. Proving the three results requires different techniques—the analysis of a classical statistical-mechanics model using Krawtchouck polynomials, graph theoretic analysis using the switching method, and analysis of the ranks of submatrices of random adjacency matrices, respectively.
@@ 1 Introduction
In our work, we consider graph states from the perspective of understanding quantum properties that lead to computational speedups. Graph states are prime candidates to study a specific quantum phenomenon, namely, entanglement, in terms of how it relates to computational complexity. On the one hand, measuring graph states in adaptive single-qubit bases allows the execution of arbitrary quantum computations through measurement-based quantum computing [75, 76]. In fact, these measurements can be restricted to the $X$-$Y$ plane of the Bloch sphere [64]. On the other hand, graph states exhibit a rich multipartite entanglement structure [49] which is in fact required for quantum speedups [28]. Understanding which properties of graph states makes them generically hard to simulate classically can therefore yield insights into the mechanisms underlying quantum speedups.
[…]
In the worst case [37] gave a tight connection between simulability and entanglement: when measured in an arbitrary $X$-$Y$ plane product basis, $d$-regular graph states are hard to simulate and highly entangled if and only if 2<d<n32<d<n-3 . But both for demonstrating quantum advantage, and when aiming to understand the intrinsic relation between hardness of simulation, multipartite entanglement, and device connectivity, it is crucial to study generic states from the family, i.e., random $d$-regular graph states [48]. [^2] Random instances from this family can be efficiently sampled [43]. This ensures that we are not drawing conclusions from isolated points in the family and makes random regular graphs a goldilocks model to study those fundamental and practical questions alike. [^3] Another, unrelated, setting in which random graph states arise naturally is measurement-based computation based on cluster states [75] with iid. erasure noise on all qubits [7], or non-deterministic gates [61]. In these scenarios there are percolation thresholds between easy-to-simulate and hard-to-simulate phases.
@@
In this paper, we therefore consider the average-case complexity of simulating uniformly random $d$-regular graph states when measured in an arbitrary $X$-$Y$ plane product basis. In doing so we study structured randomness in order to understand the relation between entanglement and complexity. From a fundamental perspective, this provides an alternative to fully random quantum states. From an experimental perspective, it provides an opportunity for demonstrating the computational capabilities of quantum processors for more structured circuits, as well as the potential for fault-tolerant implementations. Importantly, the setting we consider covers both sampling problems and universality for measurement-based quantum computing. We give evidence for the average-case complexity of these two problems in three distinct regimes of the regularity parameter, (1) the regime of $2<d=O(1)$, (2) the regime of $d\in\Theta(n^{c})$ for any $1/2<c<1$, and (3) the regime of $d\sim n/2$. To the best of our knowledge, we thus give the first ensemble of circuits which are classically intractable for any depth (in particular low depth) above a constant threshold. The best lower bound we had for such a threshold has been logarithmic depth [15, 73, 51, 84].
@@
We first give evidence for average-case complexity in the regime of $2<d=O(1)$. To this end, we first show that the output distributions of IQP circuits on random highly connected graphs constructed from $d$ random matchings measured in a random $X$-$Y$ plane basis have the anticoncentration property [2, 45] at any depth $2<d=o(n^{1/2})$. Combined with the fact that for each value of $d$ there are worst-case-hard instances in this family, this provides evidence for hardness of simulation on the same level of rigor as is known for other discrete families of IQP circuits [15, 48]. This is the first family of quantum circuits that we are aware of for which there is evidence of simulation hardness at any sublinear depth above a constant threshold, and thus a result of independent interest. We use this result to characterize the complexity of random regular graph states, noting that a random circuit from this matching model (strictly speaking a slight variation called the pairing model) at a fixed depth $d$ yields a uniformly random $d$-regular graph state with probability $O(2^{-d^{2}})$. This implies the anticoncentration property of constant-regularity graph states and thus gives evidence for their classical intractability, showing the first main result.
@@
In our second result, we show that a random $d$-regular graph with $d=\Theta(n^{c})$ for any $1/2<c<1$ has a grid graph of polynomial size as an induced subgraph (resulting from deleting some of the vertices). Since the grid graph is a universal resource for measurement-based quantum computation, this shows that an algorithm for sampling from the output distribution of those graph states in any local basis would imply a collapse to the polynomial hierarchy up to the average-case #P-hardness of a certain approximation problem (as well as being, in principle, universal resources for MBQC). To the best of our knowledge, this is the strongest known average case hardness result for any ensemble of graph states and the corresponding ensemble of IQP circuits.
@@
Finally, in our third result, we show that uniformly random graph states have geometric entanglement bounded significantly away from the maximum. This makes them nontrivial and in particular not amenable to the trivial simulation algorithm of Ref. [38], giving some evidence for their average-case complexity when measured in any local basis. We believe similar results will hold for random $cn$-regular graphs with $0<c<1/2$.
@@
Our results demonstrate that computationally universal or complex states can arise naturally from constrained randomness, and that this constraint can in fact give rise to more complexity than less structured or completely unstructured randomness. Indeed, comparing to the paradigmatic setting of quantum circuits composed of parallel Haar-random 2-qubit gates, two of our results are particularly striking. First, our result that random graph states which can be prepared in constant depth exhibit anticoncentration is provably not true for constant-depth Haar-random circuits [23, 30, 84]. These may still be average-case hard to simulate, but numerical evidence in low dimensions points against a low depth threshold for classical intractability [73]. Second, our results that uniformly random graph states are not too entangled to be useful for measurement-based computation is also provably not true for Haar-uniformly random states [38]. Thus, our results suggests that while quantum states generated by Haar-random circuits are only complex in a limited depth regime, graph states generated by random CZ circuits are complex at almost any depth.
@@ 1.1 Guide for readers
  • Anticoncentration results for IQP circuits and constant-degree regular graph states: A summary of the results and proof techniques is given in Section˜1.2.1, Section˜1.2.2, Section˜1.3.1. The results are discussed in detail and proven in Section˜3 with preliminaries in Sections˜2.1, 2.2 and 2.3.
  • Universality results for random regular graph states of intermediate degree: A summary of the results and proof techniques is given in Sections˜1.2.3 and 1.3.2. The results are proven in Section˜4.
  • Absence of a geometric entanglement barrier for random graph states of high degree: The main results and proof techniques are summarized in Sections˜1.2.4 and 1.3.3. The results are proven in Section˜5 with preliminaries in Section˜2.4.
Not to be missed in either case is the discussion in Section˜1.5.
@@ 1.2 Results
We present results on uniformly random $d$-regular graph states covering three different ranges for the regularity parameter $d$. We deal, in order, with $2<d=O(1)$, $d=\Theta(n^{c})$ for $c\in(0,1)$, and $d\sim n/2$. Along the way we will also prove several results for different random graph models, such as the pairing and matching models (described in Section˜2.2) and the uniformly random graph model. We believe these results to be of independent interest.
@@ 1.2.1 Anticoncentration of a family of random pairing and matching IQP circuits
Our first set of results deals with the average-case hardness of two families of IQP circuits of depth $d$ satisfying $d>2$ and $d=o(n^{1/2})$. These IQP circuits prepare (not always regular) graph states of degree at most $d$. Average-case hardness of constant-regularity graph states will follow from those results. To give evidence for average-case hardness of these depth-$d$ circuits we prove that their outcome distribution, when measured in a random local basis in the $X$-$Y$ plane anticoncentrates. Moreover, they contain worst-case hard instances.
@@
of the outcome distribution $p_{G,\theta}$ of the graph state on $G$ measured in $X$-$Y$-plane angles $\theta\in[0,2\pi)^{n}$. We say that the output distribution anticoncentrates if $m_{2}\in O(1)$, since in that case a constant fraction of the output probabilities must be on the order of $1/2^{n}$. Averaging over graphs $G$ and measurement angles $\theta$, the average second moment $\overline{m}_{2}=\mathbb{E}_{G,\theta}[m_{2}(G,\theta)]$ gives evidence for the $\mathsf{\#P}$-hardness of approximating the outcome probabilities of measuring states corresponding to random graphs $G$ and random angles $\theta$ up to constant relative error [48]. Intuitively, anticoncentration prohibits an average case simulator which uses trivial approximations by zero to most probabilities of most graph states in its simulation. In this sense, it gives evidence for the average-case complexity of random graph states. Assuming additional well-founded complexity theoretic assumptions, in conjunction with the worst-case hardness results of [37] this gives evidence for the hardness of sampling from such graph states to the same level of confidence as we have for other discrete families of circuits. The technical argument is reviewed in Ref. [48] and, roughly speaking, goes as follows: The circuit family we consider contains instances that are provably hard to strongly simulate up to relative error. Anticoncentration shows that most outcome probabilities have a similar order of magnitude and hence there is no detectable structure in the output distribution that would make a trivial simulation algorithm work. And finally, the ensembles we consider do not appear to have any exploitable structure that would help an algorithm designer to simulate a random instance compared to the worst-case instance, see also [14].
Specifically, we consider the following two ensembles of depth-$d$ IQP circuits, which generate graph states of degree $d$. We call the first ensemble the random pairing model. This model is obtained by choosing a uniformly random matching on $n\cdot d$ vertices and identifying $d$ vertices with a qubit. A $CZ$ gate is applied to every edge in the resulting multigraph. This model (which is well studied in graph theory [94]) is motivated by the fact that, conditioning on the graph being simple, it generates a uniformly random $d$-regular graph state. We associate (simple) graph states to the multigraphs chosen in these ways by deleting double edges and self-loops. In the second model, the random matching model, a graph state is generated by applying $CZ$ gates on $d$ independent, uniformly random matchings. In this model, for any constant $d$, a regular graph state is generated with constant probability. However, not all $d$-regular graph states admit a perfect matching [57] (and some admit many, making this distribution different from the uniformly random regular graphs, even when conditioning on simplicity). Thus, the regular graph states generated by perfect matchings are exactly those graph states with $dn/2$ edges which can be prepared in optimal depth $d$.
Figure 1: Exact values of $\mathbb{E}_{G,\theta}[m_{2}(G,\theta)]$ (evaluated via Equation˜106) for graph states drawn from (a) the random pairing ensemble $\mathcal{G}_{p}(d,n)$ and (b) the random matching ensemble $\mathcal{G}_{m}(d,n)$ for various values of $d$ as a function of $n$. The opaque (transparent) loosely dotted lines denote the asymptotic values of $m_{2}=2(3)$ of the average second moment. Solid (dashed) lines represent odd (even) values of $d$.
We show that the output distribution of both models in a random $X$-$Y$ basis anticoncentrates for any $d$ satisfying $d>2$ and $d=o(n^{1/2})$.
@@
[…]
We also show the exact value of $\mathbb{E}_{G,\theta}[m_{2}(G,\theta)]$ for different values of $n,d$ and both models in Figure˜1. We observe not only that the $d$-dependence of our results is tight, but also that the convergence to the asymptotic value occurs rapidly.
@@ 1.2.2 Anticoncentration of constant-degree regular graph states
Consider the uniform measure $\mathcal{G}_{r}(n)$ on $d$-regular graphs on $n$ vertices, and uniformly random angles $\theta$. Then, for $2<d=o(n^{1/2})$,
@@
This corollary is a straightforward consequence of the fact that conditioning on simplicity in the pairing model gives rise to uniformly random regular graphs (see Section˜2.2). Corollary˜2, together with the worst case hardness result of [37] gives evidence that approximating the outcome probabilities of regular graph states of constant degree measured random $X$-$Y$ plane bases is $\mathsf{\#P}$-hard on average. We strongly believe that random regular graph states also anticoncentrate at super-constant depth, but leave showing that to future work.
We note that some existing results about IQP circuits can be rephrased in terms of random graph state models yielding results similar to Corollary˜2 for different models of random graphs. In particular, Ref. [51] shows that IQP circuits composed of $g$ CZCZ and $Z$ gates applied uniformly at random to $n$ qubits anticoncentrate if $g=\Omega(n\log n)$ but fail to anticoncentrate for any $g=O(n)$. This directly translates to the Erdős-Rényi random graph model in which an edge is contained in the graph with probability $p=g/\binom{n}{2}$, measured in the $\pm X$ basis. Ref. [51] gives evidence that Erdős-Rényi random graphs are hard to simulate for $p\in\Omega(\log n/n)$ and suggests that they are not generically hard to simulate below that threshold. This is consistent with graph percolation, where famously Erdős and Rényi [33] showed that if $p<1/2n$, a random graphs largest connected component has size $O(\log n)$ which implies that it is efficiently simulatable using tensor-network techniques [28].